When the price of flowers increased from $5.00 to $5.70, the quantity demanded of chocolate increased from 5,550 to 6,150. What is the estimated cross-price elasticity of demand for chocolate? Round your answer to the nearest hundredth.
Answer #1

cross price elasticity of demand=(change in quantity/average quantity)/(change in price/average price)
Change in quantity=6150-5550=600
average quantity=(6150+5550)/2=5850
change in price=5.70-4.00=0.70
average price=(5.70+5.00)/2=5.35

cross price elasticity of demand=(600/5850)/(0.70/5.35)

=0.783882784

=0.78

the estimated cross-price elasticity of demand for chocolate is 0.78.